I need 100 pounds of a nut mixture that is 25% peanuts. I have mixtures that are 45% and 5% peanuts. How much of each should I use?
Answer by checkley79(3341)

.45X+.05(100-X)=.25*100
.45X+5-.05X=25
.40X=25-5
.40X=20
X=20/.40
X=50 POUNDS OF 45% PEANUT MIX.
100-50=50 POUNDS OF 5% PEANUT MIX
PROOF:
.45*50+.05*50=.25*100
22.5+2.5=25
25=25